OpenStudy (anonymous):

Suppose p varies directly with d, and p = 3 when d = 5. What is the value of d when p = 12?

OpenStudy (johnweldon1993):

So we need to find the constant here... Direct variation would be \[\large y = kx\] or for your variables \[\large p = kd\] So what we do is plug in the 2 numbers we are given...and solve for 'k' \[\large 3= 5k\] \[\large k = \frac{3}{5}\] Now that we know that...we can solve the last part...when 'p' = 12...what is D? Well \[\large p = \frac{3}{5}d\] \[\large 12 = \frac{3}{5}d\] \[\large d = 12 \times \frac{5}{3}\]